OpenStudy (anonymous):

Given the function: f(x) = 3*x + 3. Find the value of y when x = -8.

OpenStudy (anonymous):

do i put -8 in both spots?

OpenStudy (anonymous):

no here f(x) is y so put x=-8 in right hand side only

OpenStudy (anonymous):

y=3*-8+3?

OpenStudy (anonymous):

yes can you tell me final value?

OpenStudy (anonymous):

y=-24+3? is that right

OpenStudy (anonymous):

final y=-24+3=-21

OpenStudy (anonymous):

wait so is -21 the only thing i put?

OpenStudy (anonymous):

yes y=-21 when x = -8 for equation y=f(x) = 3*x + 3.

OpenStudy (anonymous):

ok thank u

OpenStudy (anonymous):

you are welcome.